# A deep learning approach for predicting multiple observables in Au+Au collisions at RHIC

### v1

@@ Abstract
We develop a neural network model, based on the processes of high-energy heavy-ion collisions, to study and predict several experimental observables in Au+Au collisions. We present a data-driven deep learning framework for predicting multiple bulk observables in Au+Au collisions at RHIC energies. A single neural network is trained exclusively on experimental measurements of charged-particle pseudorapidity density distributions, transverse-momentum spectra and elliptic flow coefficients over a broad range of collision energies and centralities. The network architecture is inspired by the stages of a heavy-ion collision – from the quark-gluon plasma to chemical and kinetic freeze-out – and employs locally connected hidden layers and a structured input design that encode basic geometric and kinematic features of the system. We demonstrate that these physics-motivated choices significantly improve test performance compared to purely fully connected baselines. The trained model is then used to predict the above observables at collision energies not yet explored experimentaly at RHIC, and the results are validated using the energy dependence of the total charged-particle multiplicity per participant pair as well as comparisons to a CLVisc hydrodynamic calculation with TRENTo initial conditions. Our findings indicate that such physics-guided neural networks can serve as efficient surrogates to fill critical data gaps at RHIC and to support further phenomenological studies of QGP properties.
@@ I Introduction
Due to advanced technology, the development of GPUs and the abundance of big data, machine learning has been developed considerably in the last few years and has been increasingly applied to scientific researches including high-energy nuclear physics He et al. (2023); de Oliveira et al. (2017); Steinheimer et al. (2019); Huang et al. (2022); Mallick et al. (2023); Omana Kuttan et al. (2020); Pang et al. (2019); Mengel et al. (2023); Guo et al. (2023); Sun et al. (2024); Paganini et al. (2018); Sun et al. (2025); Gao et al. (2021); Cao et al. (2023); Mumpower et al. (2022). In Ref. de Oliveira et al. (2017), a novel Generative Adversarial Networks architecture (GAN) is applied to the production of jet images. In Ref. Steinheimer et al. (2019), a state-of-the-art machine learning method is adopted to study the coordinate and momentum space configurations of the net baryon number in heavy ion collisions that undergo spinodal decomposition, due to a first-order phase transition. In Ref. Huang et al. (2022), a point cloud network with dynamical edge convolution is employed to identify events with critical fluctuations through supervised learning, and pick out a large fraction of signal particles used for decision-making in each single event. In Ref. Mallick et al. (2023), a deep learning feed-forward network is used for estimating elliptic flow coefficients in heavy-ion collisions. In Ref. Omana Kuttan et al. (2020), a deep learning-based event characterization method is proposed for fast, online event-by-event impact parameter determination at the CBM experiment. These studies demonstrate that machine learning has become a powerful tool for studies in high-energy nuclear physics.
@@ Acknowledgements.
We thank for the helpful discussions with Prof. Ze-Fang Jiang, Prof. Wei Dai and Prof. Dirk H. Rischke. The work is supported by the China Scholarship Council (CSC) (Jun-Qi Tao), the CUHK-Shenzhen University development fund under grant No. UDF01003041 and UDF03003041, the Shenzhen Peacock fund under No. 2023TC0179 and the Ministry of Science and the Technology of China under Grant No. 2024YFA1611004 (Kai Zhou), the National Natural Science Foundation of China (NSFC) under Grants No. 12535010 (Ben-Wei Zhang).

### v2

@@ Abstract
We develop a neural network model, based on the processes of high-energy heavy-ion collisions, to study and predict several experimental observables in Au+Au collisions. We present a data-driven deep learning framework for predicting multiple bulk observables in Au+Au collisions at RHIC energies. A single neural network is trained exclusively on experimental measurements of charged-particle pseudorapidity density distributions, transverse-momentum spectra and elliptic flow coefficients over a broad range of collision energies and centralities. The network architecture is inspired by the stages of a heavy-ion collision, from the quark-gluon plasma to chemical and kinetic freeze-out, and employs locally connected hidden layers and a structured input design that encodes basic geometric and kinematic features of the system. We demonstrate that these physics-motivated choices significantly improve test performance compared to purely fully connected baselines. The trained model is then used to predict the above observables at collision energies not yet explored experimentally at RHIC, and the results are validated using the energy dependence of the total charged-particle multiplicity per participant pair as well as comparisons to a CLVisc hydrodynamic calculation with TRENTo initial conditions. Our findings indicate that such physics-guided neural networks can serve as efficient surrogates to fill critical data gaps at RHIC and to support further phenomenological studies of QGP properties.
@@ I Introduction
Due to advanced technology, the development of GPUs and the abundance of big data, machine learning has been developed considerably in the last few years and has been increasingly applied to scientific researches including high-energy nuclear physics Zhou et al. (2024); He et al. (2023); de Oliveira et al. (2017); Steinheimer et al. (2019); Huang et al. (2022); Mallick et al. (2023); Omana Kuttan et al. (2020); Pang et al. (2019); Mengel et al. (2023); Guo et al. (2023); Sun et al. (2024); Paganini et al. (2018); Sun et al. (2025); Gao et al. (2021); Cao et al. (2023); Mumpower et al. (2022); Du et al. (2020); Kuttan et al. (2021); Zhao et al. (2022); Omana Kuttan et al. (2023). In Ref. de Oliveira et al. (2017), a novel Generative Adversarial Networks architecture (GAN) is applied to the production of jet images. In Ref. Steinheimer et al. (2019), a state-of-the-art machine learning method is adopted to study the coordinate and momentum space configurations of the net baryon number in heavy ion collisions that undergo spinodal decomposition, due to a first-order phase transition. In Ref. Huang et al. (2022), a point cloud network with dynamical edge convolution is employed to identify events with critical fluctuations through supervised learning, and pick out a large fraction of signal particles used for decision-making in each single event. In Ref. Mallick et al. (2023), a deep learning feed-forward network is used for estimating elliptic flow coefficients in heavy-ion collisions. In Ref. Omana Kuttan et al. (2020), a deep learning-based event characterization method is proposed for fast, online event-by-event impact parameter determination at the CBM experiment. These studies demonstrate that machine learning has become a powerful tool for studies in high-energy nuclear physics.
@@ Acknowledgements.
We thank for the helpful discussions with Prof. Wei Dai, Prof. Ze-Fang Jiang and Prof. Dirk H. Rischke. The work is supported by the China Scholarship Council (CSC) (Jun-Qi Tao), the CUHK-Shenzhen University development fund under grant No. UDF01003041 and UDF03003041, the Shenzhen Peacock fund under No. 2023TC0179 and the Ministry of Science and the Technology of China under Grant No. 2024YFA1611004 (Kai Zhou), the National Natural Science Foundation of China (NSFC) under Grants No. 12535010 (Ben-Wei Zhang).